Circular No. 9016

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 B2 (LINEAR) An apparently asteroidal object that was discovered by the LINEAR

survey (discovery observation tabulated below) has been found to show cometary appearance by CCD astrometrists elsewhere, following posting on the Minor Planet Center’s ‘NEOCP’ webpage. F. Hormuth (Calar Alto 1.23-m reflector, Jan. 30.2 UT) writes that an unfiltered, stacked, 600-s image shows a coma diameter of ∼ 5′′ and a short, faint tail of length ∼ 10′′ in p.a. ∼ 340^{o}. J. Young (Table Mountain 0.61-m f/16 Cassegrain reflector, Jan. 31.50–31.55) notes a coma of diameter 8′′ with a small bright area slightly southeast of a broad and very short extension spanning p.a. 310^{o}- 340^{o}. R. S. McMillan (0.9-m f/3 Spacewatch reflector, Jan. 31.5; apparent independent discovery) reports a short 8′′ tail toward the north.

2009 UT α_{2000} δ_{2000} Mag. Observer
Jan. 29.47388 14^{h}50^{m}27.^{s}51 −15^{o}38′30′′.9 19.4 LINEAR

The available astrometry, the following preliminary parabolic orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2009-B91.

T = 2009 Mar. 10.773 TT  ω = 193.^{o}959
Ω = 18.905 }2000.0
q = 2.32776 AU  i = 156.994

COMET P/2008 Y1 (BOATTINI) Additional astrometry has shown that this comet (cf. IAUC 9007) is of

short period, the following orbital elements taken from MPEC 2009-B62:

T = 2009 Feb. 25.1897 TT  ω = 162.^{o}4578
e = 0.732532  Ω = 259.7532 }2000.0
q = 1.270613 AU  i = 8.8036
a = 4.750529 AU n^{o} = 0.0951900 P = 10.35 years

COMET P/2008 Y3 (McNAUGHT) Additional astrometry has shown that this comet (cf. IAUC 9009) is of

short period; the following orbital elements were published on MPC 64767: Epoch = 2009 Jan. 9.0 TT, T = 2009 Jan. 11.9366 TT, e = 0.447539, q = 4.434228 AU, ω = 238^{o}.2745, Ω = 262^{o}.9356, i = 38^{o}.8137 (equinox 2000.0), P = 22.74 years.

2009 January 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green